Pls. help me solve this problem.
The toal of two investments is 25,000. A Part of it was invested at 7 percent and one at 9 percent. The annual interest from the 7 percent investment is $470 more than the 9 percent investment. How much is invested at each rate?

Answer by mananth(16946)

The toal of two investments is 25,000. A Part of it was invested at 7 percent and one at 9 percent. The annual interest from the 7 percent investment is $470 more than the 9 percent investment. How much is invested at each rate?
A Part 7 percent------------$x
Part B 9 percent-----------$y
x+y=25000..........(1)
Interest from 9 % = 0.09y
interest from 7% = 0.07x
0.07x-470=0.09y
multiply by 100
7x-47000=9y
7x-9y=47000......(2)
multiply (1) by -7
we get -7x-7y=-175000
add to (2)
-16y=-175000+47000
-16y=-128000
/-16
y= $8000 ===> investment at 9%
Balance $17,000 at 7%